OpenStudy (anonymous):

I have couple of questions to ask:) please HELP! 1. lim x->0 {x-xcosx}/{sin^2(3x)} 2. lim x->1 cos^-1 (lnx) 3. lim x->0 cos {pi/sqrt(19-3sec2x)} 4. lim x->0+ [{1/x^(1/3)} - {1/(x-1)^(4/3)}]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{x-x\cos x}{\sin^23x}\] Quite a few ways to go about this one. Keep in mind the special trig limits: \[\lim_{x\to0}\frac{\sin ax}{ax}=1\\ \lim_{x\to0}\frac{1-\cos ax}{ax}=0\] So for this first limit, we're going to have to do some rewriting to make it somewhat resemble these special ones. \[\begin{align*}\lim_{x\to0}\frac{x-\cos x}{\sin^23x}&=\lim_{x\to0}\frac{x}{\sin3x}\cdot\frac{1}{\sin 3x}\cdot\frac{1-\cos x}{1}\\ &=\lim_{x\to0}\frac{x}{\sin3x}\cdot\lim_{x\to0}\frac{1}{\sin3x}\cdot\lim_{x\to0}\frac{1-\cos x}{1}\\ &=\left(\lim_{x\to0}\frac{x}{\sin3x}\cdot\frac{3}{3}\right)\left(\lim_{x\to0}\frac{1}{\sin3x}\cdot\frac{3x}{3x}\right)\left(\lim_{x\to0}\frac{1-\cos x}{1}\cdot\frac{x}{x}\right)\\ &=\color{red}{\lim_{x\to0}\frac{x}{3(3x)}}\cdot\left(\lim_{x\to0}\frac{3x}{\sin3x}\right)^2\cdot\lim_{x\to0}\frac{1-\cos x}{x} \end{align*}\] The red part is the "residue" from multiplying by 1. Should be a piece of cake from here.

OpenStudy (anonymous):

The second limit is easier. Directly substituting should do it. \[\ln1=0~~\Rightarrow~~\cos^{-1}0=\cdots\]

OpenStudy (anonymous):

Likewise for the third limit. \[\sec2(0)=\sec0=1\] so \[\cos\left(\frac{\pi}{\sqrt{19-3}}\right)=\cos\frac{\pi}{4}=\cdots\]